# Approximating the coefficients of the Bessel functions

### v4

@@ 9. Asymptotics of the coefficients of Bessel functions
In this section, we determine the asymptotics of the coefficients of Bessel functions in each of the regimes mentioned in Subsection˜1.2. Note that we determine the asymptotics for the coefficients of the terms with a fixed degree and the analyses for different degrees are separate. Therefore, additional assumptions are required to deduce the convergence of the Bessel functions themselves rather than only their coefficients. For more discussion about this direction, see Subsection˜10.3.
@@ 9.2. Coefficients for the $BC^{N}$ root system
    The diagonal entry $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\lambda})^{-1}$.
@@
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]_{\lambda\nu}^{-1}=0$.
@@
$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{BC}([r_{l}])(2c_{0},2c_{1})\right)^{-1}\times$
@@ 9.3. Coefficients for the $D^{N}$ root system
Lemma 9.9.
Suppose $|\theta N|\rightarrow\infty$. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{M}^{D}[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{M}^{D}[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{M}^{D}[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.10. ∎
Lemma 9.10.
Suppose $|\theta N|\rightarrow\infty$. Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$. The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
$(\theta N)^{-k+\ell(\lambda)}N^{-\ell(\nu)}\prod_{j=1}^{N}(1+2(j-1)\theta)^{-1}(\mathcal{M}^{D}[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
Proof.
See Theorem˜8.16. ∎
Corollary 9.11.
If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then
$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}(\theta N)^{-k}\theta^{\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}2^{r_{l}-1}r_{l}\right)^{-1}\times$
$\left(1+o_{N}(1)+(1+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\right)r(a)r(x)$
$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}(\theta N)^{-k}(\theta N)^{\min(\ell(r),\ell(s))}N^{-\max(\ell(r),\ell(s))}\times$
$\bigg[\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1)$
$+(\mathcal{M}^{D}[k]^{-1}_{rs}+\mathcal{M}^{D}[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
Proof.
See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10. Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$. Furthermore, we obtain the term $\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$ in $J_{a}^{D^{N}(\theta)}$ from $e,J_{a}^{D^{N}(\theta)}]_{D^{N}(\theta)}=e(a)$. ∎
Lemma 9.12.
Suppose $\theta N\rightarrow c\in\mathbb{C}$ and $2c$ is not a negative integer. Suppose $k\geq 1$ and define $M\triangleq([p_{\lambda},p_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}(\mathcal{W}^{A}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{A}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{A}(2c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{A}(2c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.12. ∎
Lemma 9.13.
Suppose $\theta N\rightarrow c\in\mathbb{C}$, $2c$ is not a negative integer, and assume that $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large. Suppose $k\geq 1$ and define $M\triangleq([ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)})_{\lambda,\nu\in\Gamma_{\text{even}}[k]}$. Suppose $\lambda,\nu\in\Gamma_{\text{even}}[k]$ and $\ell(\lambda)\leq\ell(\nu)$.
  1. (a)
    The entries $M^{-1}_{\lambda\nu}$ and $M^{-1}_{\nu\lambda}$ equal
    $N^{-\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)^{-1}(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{\lambda\nu}+o_{N}(1)).$
  2. (b)
    The diagonal entry $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\lambda})^{-1}$.
  3. (c)
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]_{\lambda\nu}^{-1}=0$.
Proof.
See Theorem˜8.23. ∎
Corollary 9.14.
If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then
$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
$J^{D^{N}(\theta)}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}\right)^{-1}\bigg[\prod_{l=1}^{\ell(r)}W^{A}([r_{l}])(2c)^{-1}+o_{N}(1)$
$+\left(\prod_{l=1}^{\ell(r)}W^{D;\,\text{odd}}(\{1\}\cup\{2,\ldots,r_{l}+1\})(2c)^{-1}+o_{N}(1)\right)\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)r(x)$
$+\sum_{r,s\in\Gamma_{\text{even}}[k],\,r\not=s}N^{-\max(\ell(r),\ell(s))}\bigg[\mathcal{W}^{A}(2c)[k]^{-1}_{rs}+\mathcal{W}^{A}(2c)[k]^{-1}_{sr}+o_{N}(1)$
$+(\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{rs}+\mathcal{W}^{D;\,\text{odd}}(2c)[k]^{-1}_{sr}+o_{N}(1))\frac{e(a)e(x)}{\prod_{j=1}^{N-k}(1+2(j-1)\theta)}\bigg]r(a)s(x).$
Proof.
@@ 10. Applications
In this section, we discuss applications of the main results of this paper. Some settings that we consider include the weak convergence of the measures in Conjecture˜1.6 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
@@ 10.1. Weak convergence to the free convolution
The results of this subsection are based on the assumption of Conjecture˜1.6. Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems. In particular, we prove Corollary˜1.7. For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.8.
@@
Assume that $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
@@
Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.4. Then, we have that
@@
which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.4. Thus, $u$ is in the support of $\epsilon$.
However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$. Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.4, the result follows. ∎
In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$. Afterwards, it is straightforward to deduce Corollaries˜1.7 and 1.8.
@@
    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{d}^{a}$ and $m_{d}^{b}$ are real numbers such that
    $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}b(N)_{i}^{d}}{(\theta N)^{d}N}=m_{d}^{b}$
    for $d\geq 1$. Let $a_{(k)}$ and $b_{(k)}$ for $k\geq 1$ be real numbers that solve
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{d}\right]}{(\theta N)^{d}N}=\sum_{\pi\in NC(d)}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta_{0}N)^{2d}N}$
    $=\sum_{\pi\in NC^{\text{even}}(2d)}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}c_{i}^{2d}\right]}{(\theta N)^{2d}N}=\sum_{\pi\in NC^{\text{even}}(2d)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.14. Thus, part (D) of Lemma˜10.15 implies that for all $\lambda\in\Gamma$,
@@
(B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.22, respectively, and follow the same idea as the proof of (A). ∎
@@
(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition. Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$. It follows from [billingsley]*Theorem 30.2 that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in distribution to $\mu$.
(B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$. It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero. Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$. Afterwards, we can follow the steps of (A). ∎
@@
[…]
Suppose $d\geq 1$ and $\pi\in NC^{\text{even}}(2d)$. Recall that $o(\pi)$ is the number of $i\in[2d]$ such that $i$ is not the minimal element of a block of $\pi$ and $d(i;\,\pi)$ is odd. Suppose $\pi=B_{1}\sqcup\cdots\sqcup B_{m}$ such that $\min(B_{j})<\min(B_{j+1})$ for $1\leq j\leq m-1$. Suppose $i\in B_{j}$ for $j\in[m]$. We have that
@@
the equation also holds after replacing $a_{(|B|)}$ with $b_{(|B|)}$. By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\\ \left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. ∎

### v5

@@ 9. Asymptotics of the coefficients of Bessel functions
In this section, we determine the asymptotics of the coefficients of Bessel functions in each of the regimes mentioned in Subsection˜1.4. Note that we determine the asymptotics for the coefficients of the terms with a fixed degree and the analyses for different degrees are separate. Therefore, additional assumptions are required to deduce the convergence of the Bessel functions themselves rather than only their coefficients. For more discussion about this direction, see Subsection˜10.3.
@@ 9.2. Coefficients for the $BC^{N}$ root system
    The diagonal entry $\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\lambda}^{-1}$ equals $(\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\lambda})^{-1}$.
@@
    If $\ell(\lambda)=\ell(\nu)$ and $\lambda\not=\nu$, then $\mathcal{W}^{BC}(c_{0},c_{1})[k]_{\lambda\nu}^{-1}=0$.
@@
$J^{BC^{N}(\theta_{0},\theta_{1})}_{a}[k](x)=\sum_{r\in\Gamma_{\text{even}}[k]}N^{-\ell(r)}\left(\pi(r)\prod_{l=1}^{\ell(r)}r_{l}W^{BC}([r_{l}])(c_{0},c_{1})\right)^{-1}\times$
@@ 9.3. Coefficients for the $D^{N}$ root system
The framework of this subsection is analogous to that of the $A^{N-1}$ root system. By [demni_type_d], we have that
(20) $J_{a}^{D^{N}(\theta)}(x)=J_{a}^{BC^{N}(\theta,0)}(x)+\frac{e(a)e(x)}{\prod_{i=1}^{N}(1+2(i-1)\theta)}J_{a}^{BC^{N}(\theta,1)}(x).$
Therefore, it suffices to consider the coefficients of the Bessel function for the $BC^{N}$ root system.
@@ 10. Applications
In this section, we discuss applications of the main results of this paper. Some settings that we consider include the weak convergence of the measures in Conjecture˜1.5 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
@@ 10.1. Weak convergence to the free convolution
The results of this subsection are based on the assumption of Conjecture˜1.5. Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems. In particular, we prove Corollary˜1.7. For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.8.
@@
Assume that $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
@@
Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.3. Then, we have that
@@
which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.3. Thus, $u$ is in the support of $\epsilon$.
However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$. Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.3, the result follows. ∎
In the following corollary, we prove the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of power sums after assuming that $a$ and $b$ converge in terms of power sums as $N\rightarrow\infty$. Afterwards, it is straightforward to deduce Corollaries˜1.7 and 1.8.
@@
    Suppose $\lim_{N\rightarrow\infty}\theta N=\infty$ and that $m_{d}^{a}$ and $m_{d}^{b}$ are real numbers such that for all $d\geq 1$,
    $\lim_{N\rightarrow\infty}\frac{p_{(d)}(a(N))}{(\theta N)^{d}N}=m_{d}^{a}\text{ and }\lim_{N\rightarrow\infty}\frac{p_{(d)}(b(N))}{(\theta N)^{d}N^{1}}=m_{d}^{b}.$
    Let $a_{(k)}$ and $b_{(k)}$ for $k\geq 1$ be real numbers that solve
@@
    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[p_{\lambda}(c)\right]}{(\theta N)^{|\lambda|}N^{\ell(\lambda)}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\left[p_{\lambda}(c)\right]}{(\theta_{0}N)^{|\lambda|}N^{\ell(\lambda)}}$
    $=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\text{even}}(\lambda_{i})}(1+c)^{o(\pi)}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
    for $d\geq 1$. Assume that for $N\geq 2$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.5. Then, for all $\lambda\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{\operatorname{\mathbb{E}}_{c\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[p_{\lambda}(c)\right]}{(\theta N)^{|\lambda|}N^{\ell(\lambda)}}=\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\text{even}}(\lambda_{i})}\prod_{B\in\pi}2^{|B|-1}|B|(a_{(|B|)}+b_{(|B|)}).$
@@
For $N\geq 2$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15. Thus, part (D) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
@@
(B) and (C): For (B) and (C), we can use Corollaries˜7.5 and 8.10, respectively, and follow the same idea as the proof of (A). ∎
@@
(A): By part (A) of Corollary˜10.2, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}(x\mapsto\frac{x}{\theta N})$ converges in terms of power sums to $\mu$ as $N\rightarrow\infty$.
(B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$. It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero. Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 2}$ converges in terms of power sums to $\tilde{\mu}$. ∎
@@
We use part (B) of Corollary˜10.2. Suppose $d\geq 1$ and $\pi\in NC^{\text{even}}(2d)$. Recall that $o(\pi)$ is the number of $i\in[2d]$ such that $i$ is not the minimal element of a block of $\pi$ and $d(i;\,\pi)$ is odd. Suppose $\pi=B_{1}\sqcup\cdots\sqcup B_{m}$ such that $\min(B_{j})<\min(B_{j+1})$ for $1\leq j\leq m-1$. Suppose $i\in B_{j}$ for $j\in[m]$. We have that
@@
the equation also holds after replacing $a_{(|B|)}$ with $b_{(|B|)}$. By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto\frac{x}{\theta_{0}N})+\frac{1}{2}\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}(x\mapsto-\frac{x}{\theta_{0}N})$ converges in terms of power sums to $\mu$. ∎